OpenStudy (anonymous):

Find the equation of the line that passes through (4, 5) with a slope of 2. Write your answer in Slope-Intercept Form. y = 2x + 1 y = 2x - 5 y = 2x -1 y = 2x - 3

OpenStudy (luigi0210):

Do you know the point-slope formula?

OpenStudy (anonymous):

No Im sorry thats all the information given:(

OpenStudy (luigi0210):

Point-slope formula: \[\LARGE y-y_{1}=m(x-x_{1})\] \[\Large (x_{1},~ y_{1})\] \[\LARGE m=SLOPE\]

OpenStudy (anonymous):

Thanks!

OpenStudy (uri):

 So, y-5=2(x-4)